# Fermionic Averaged Circuit Eigenvalue Sampling

### v1

@@ V Conclusion
Ultimately, we would like to utilize these protocols to not only learn circuit-level noise, but to also mitigate it. With this in mind, a important feature of our protocol is that the design matrix is necessarily block diagonal. In the aforementioned question of gauge parameters, it is already clear that each unlearnable function thus depends only on gate eigenvalues from the same Majorana-degree sector. Additionally, noisy expectation values of fixed Majorana-degree observables are necessarily linear functions of the ideal expectation values, even in the setting where these observables are not Majorana monomials (e.g. measuring the expected energy for a Hamiltonian of fixed Majorana degree). This contrasts with the setting of Pauli observables under Clifford-circuit evolution, and is reminiscent of training methods for error mitigation [28, 29]. We leave a careful consideration of this problem for future work.
Acknowledgements.
@@ Appendix A Pseudocode Algorithms
In this section we provide pseudocode implementations for the steps of our protocol described in fig. 1. Algorithm 1 is the main function, which learns the gate eigenvalues $\boldsymbol{\xi}$ describing an error model $\mathcal{M}$ for a collection of circuits $\mathcal{C}$ in a given number of measurements per circuit $S$. First, the design matrix $\mathbf{A}$ is computed using algorithm 2. The rows of this matrix are indexed by circuits in $\mathcal{C}$, and the columns are indexed by noise parameters, such that $A_{jk}$ corresponds to the number of times the noise parameter $k$ appears in circuit $j$. In algorithm 3, a probability distribution of outcomes is generated for each circuit according to its type (either “$x$” or “$z$”). These distributions are then input to algorithm 4, where they are used to calculate the circuit eigenvalues by eq. 48 or eq.s (49) and (50). Finally, this is used to produce the gate eigenvalues $\boldsymbol{\xi}$ using a similar linear inversion procedure to the ACES protocol. In this procedure, $\mathbf{A}$ is truncated to only include its elements which lie above some given cutoff $\delta$. If $\mathbf{A}$ is still full rank, we solve the linear equation $\mathbf{A}\mathbf{x}=\mathbf{b}$ relating the logarithms of the circuit, $\mathbf{b}$, and gate eigenvalues $\mathbf{x}$. Finally, we recover the gate eigenvalues themselves by exponentiating the components of $\mathbf{x}$.
@@
3:$\mathbf{A}\leftarrow\textsc{Design}(\mathcal{M})$
4:$\boldsymbol{\Lambda}\leftarrow[\ \ ]$triangleright Initialize an empty list
@@
12:$\hat{\boldsymbol{\xi}}\leftarrow\textsc{GateEigs}({\boldsymbol{\Lambda}},\mathbf{A})$
13:$\hat{\boldsymbol{\xi}}$
@@
$\mathcal{M}$
2:$J\leftarrow|\mathcal{C}|$
3:$K\leftarrow|\mathcal{G}|$
4:$\mathbf{A}\leftarrow\mathbf{0}_{J\times K}$triangleright Initialize $\mathbf{A}$ as an all-zeros matrix.
@@
14:end for  # Check if $A$ has full column-rank.
@@
16:     throw(“$\mathbf{A}$ is not full rank.
@@
5:     $\mathcal{U}^{\prime}\leftarrow\textsc{Twirl}(\mathcal{U})$triangleright twirled version of $\mathcal{U}$
@@
1:$\boldsymbol{p}$
2:# $\mathbf{M}$ is calculated using lemma 4.
@@
6$\triangleright$ else type($\boldsymbol{p}$) = $x$
7:# Apply eq. 49 and eq. 50
@@
1:$\mathbf{A},\hat{\boldsymbol{\Lambda}},\delta$
2:$J,K\leftarrow\textsc{size}(\mathbf{A})$
@@
5:$\mathbf{A}\leftarrow\mathbf{A}[\hat{\boldsymbol{\Lambda}}>\delta,\,:\,]$
@@
8:     throw(“$\mathbf{A}$ is not full rank with cutoff $\delta$.”)
@@
10:$\hat{\mathbf{b}}\leftarrow-\log(\hat{\boldsymbol{\Lambda}})$
11:$\hat{\mathbf{x}}\leftarrow\textsc{solve}(\mathbf{A},\hat{\mathbf{b}})$triangleright Solve $\mathbf{A}\hat{\mathbf{x}}=\hat{\mathbf{b}}$
12:$\hat{\boldsymbol{\xi}}\leftarrow\exp(-\hat{\mathbf{x}})$
13:$\hat{\boldsymbol{\xi}}$
@@ B.1 Proof of lemma 4
eq. 81 immediately follows from the definition eq. 9 and Vieta’s formulas.
Now consider eq. 82, and let $\boldsymbol{b}\coloneqq-\mathbf{1}_{j}\oplus\mathbf{1}_{\ell-j}$ from the definition in eq. 9. We have
@@
yielding the result eq. 82.
To show eq. 83, we expand the polynomial in the right side of eq. 81 by the binomial theorem
@@
The limits of the sum are such that the binomial coefficients are well-defined (this is redundant since our convention is such that the binomial coefficients are zero if the arguments are outside of the bounds). The lower limit is such that $m\geq 0$ to satisfy the lower bound of $\binom{j}{m}$ and $k-m\leq\ell-j$ to satisfy the upper bound of $\binom{\ell-j}{k-m}$. The upper bound is such that $m\leq j$ to satisfy the upper bound of $\binom{j}{m}$ and $m\leq k$ to satisfy the lower bound of $\binom{\ell-j}{k-m}$. Equation 89, together with eq. 81, gives the result.
To show eq. 84, let $v_{j}(u)\equiv u^{j}$ for $j\in\{0\}\cup[\ell]$. We have, again by eq. 9 and Vieta’s formulas,
@@ B.2 Proof of lemma 6
Since exchanging rows $j$ and $\ell-j$ in $\mathbf{M}^{(\ell)}$ is equivalent to multiplying the argument of the symmetric polynomial in eq. 9 by a factor of $-1$, we have
@@
where the second line follows from the identity $e_{k}(-\boldsymbol{v})=(-1)^{k}e_{k}(\boldsymbol{v})$. To show the result for the other ordering, we utilize the reciprocal polynomial of $p(x)$, given by $p^{*}(x)=x^{d}p(1/x)$, where $d$ is the degree of $p$. In particular, the degree-$k$ coefficient of $p$ is the degree-$(d-k)$ coefficient of $p^{*}$. For any $k$, we have by lemma 4
@@
Together, eq. 100 and eq. 105 complete the proof. ∎
B.3 Proof of lemma 7
@@
Each factor in parentheses is a FLO transformation, since $\gamma_{j}\gamma_{k}=\exp\left(\frac{\pi}{2}\gamma_{j}\gamma_{k}\right)$, and $\gamma_{1}\in\mathds{F}(2n)$ by the definition below eq. 3. Thus, we can perform the FLO twirl on $\mathcal{E}$ by first performing a Pauli twirl, mapping $\mathcal{E}\rightarrow\mathcal{E}^{\mathds{P}(n)}$, and then twirling over the full group of FLO unitary transformation, mapping $\mathcal{E}^{\mathds{P}(n)}\rightarrow(\mathcal{E}^{\mathds{P}(n)})^{\mathds{F}(2n)}=\mathcal{E}^{\mathds{F}(2n)}$.
By applying the Jordan-Wigner transform to eq. 22, together with the first statement of the lemma, we have
[…]
@@
[…]
@@
[…]
@@
[…]
@@ B.4 Proof of lemma 11
For an ensemble of noisy FLO circuits $\mathcal{C}_{r}$,
@@
We begin with the statement for $\mathcal{C}_{z}$. We have
@@
Applying the argument leading to eq. 37 gives
@@
Circuits in $\mathcal{C}_{z}$ are chosen to interact as the identity on net. This gives
@@
$\mathrm{P}^{(j)}_{0,\ell}=2^{-n}\mathbf{d}^{(n)}\cdot\mathbf{M}^{(n)}\cdot\boldsymbol{\Lambda}_{\mathrm{even}}^{(j)}.$ (133)
In the $\mathcal{C}_{x}$-case, instead of eq. 122, we have
@@
The corresponding probability for the “--” outcome on the first bit is given by
@@
$\frac{1}{2}(\mathrm{P}^{(j)}_{+,\ell}+\mathrm{P}^{(j)}_{-,\ell})$ $=2^{-n}\mathbf{d}^{(n-1)}\cdot\mathbf{M}^{(n-1)}\cdot\boldsymbol{\Lambda}_{\mathrm{even}\setminus\{2n\}}^{(j)}$ (148)
$\frac{1}{2}(\mathrm{P}^{(j)}_{+,\ell}-\mathrm{P}^{(j)}_{-,\ell})$ $=2^{-n}\mathbf{d}^{(n-1)}\cdot\mathbf{M}^{(n-1)}\cdot\boldsymbol{\Lambda}_{\mathrm{odd}}^{(j)}$ (149)

### v2

@@ 5 Conclusion
Ultimately, we would like to utilize these protocols to not only learn circuit-level noise, but to also mitigate it. With this in mind, an important feature of our protocol is that the design matrix is necessarily block diagonal. In the aforementioned question of gauge parameters, it is already clear that each unlearnable function thus depends only on gate eigenvalues from the same Majorana-degree sector. Additionally, noisy expectation values of fixed Majorana-degree observables are necessarily linear functions of the ideal expectation values, even in the setting where these observables are not Majorana monomials (e.g. measuring the expected energy for a Hamiltonian of fixed Majorana degree). This contrasts with the setting of Pauli observables under Clifford-circuit evolution, and is reminiscent of training methods for error mitigation [30, 31]. We leave a careful consideration of this problem and other applications of our results to quantum simulation for future work.
@@ Appendix A Pseudocode Algorithms
In this section we provide pseudocode implementations for the steps of our protocol described in fig.˜1. Algorithm 1 is the main function, which learns the gate eigenvalues $\boldsymbol{\xi}$ describing an error model $\mathcal{M}$ for a collection of circuits $\mathcal{C}$ in a given number of measurements per circuit $S$. First, the design matrix $\mathbf{A}$ is computed using algorithm˜2. The rows of this matrix are indexed by circuits in $\mathcal{C}$, and the columns are indexed by noise parameters, such that $A_{jk}$ corresponds to the number of times the noise parameter $k$ appears in circuit $j$. In algorithm˜3, a probability distribution of outcomes is generated for each circuit according to its type (either “xx” or “zz”). These distributions are then input to algorithm˜4, where they are used to calculate the circuit eigenvalues by eq.˜48 or eq.s (49) and (50). Finally, this is used to produce the gate eigenvalues $\boldsymbol{\xi}$ using a similar linear inversion procedure to the ACES protocol. In this procedure, $\mathbf{A}$ is truncated to only include its elements which lie above some given cutoff $\delta$. If $\mathbf{A}$ is still full rank, we solve the linear equation $\mathbf{A}\mathbf{x}=\mathbf{b}$ relating the logarithms of the circuit, $\mathbf{b}$, and gate eigenvalues $\mathbf{x}$. Finally, we recover the gate eigenvalues themselves by exponentiating the components of $\mathbf{x}$.
@@
[…]
@@
[…]
@@
[…]
@@
14:end for  # Check if AA has full column-rank.
@@
16:  throw($\mathbf{A}$ is not full rank.
@@
5:  $\mathcal{U}^{\prime}\leftarrow\textsc{Twirl}(\mathcal{U})$ $\triangleright$ twirled version of $\mathcal{U}$
@@
[…]
2:# $\mathbf{M}$ is calculated using lemma˜4.
@@
[…]
7:# Apply eq.˜49 and eq.˜50
@@
[…]
@@
[…]
@@
8:  throw($\mathbf{A}$ is not full rank with cutoff $\delta$.”)
@@
[…]
@@ B.1 Proof of lemma˜4
eq.˜81 immediately follows from the definition eq.˜9 and Vieta’s formulas.
Now consider eq.˜82, and let $\boldsymbol{b}\coloneqq-\mathbf{1}_{j}\oplus\mathbf{1}_{\ell-j}$ from the definition in eq.˜9. We have
@@
yielding the result eq.˜82.
To show eq.˜83, we expand the polynomial in the right side of eq.˜81 by the binomial theorem
@@
The limits of the sum are such that the binomial coefficients are well-defined (this is redundant since our convention is such that the binomial coefficients are zero if the arguments are outside of the bounds). The lower limit is such that $m\geq 0$ to satisfy the lower bound of $\binom{j}{m}$ and $k-m\leq\ell-j$ to satisfy the upper bound of $\binom{\ell-j}{k-m}$. The upper bound is such that $m\leq j$ to satisfy the upper bound of $\binom{j}{m}$ and $m\leq k$ to satisfy the lower bound of $\binom{\ell-j}{k-m}$. Equation˜89, together with eq.˜81, gives the result.
To show eq.˜84, let $v_{j}(u)\equiv u^{j}$ for $j\in\{0\}\cup[\ell]$. We have, again by eq.˜9 and Vieta’s formulas,
@@ B.2 Proof of lemma˜6
Since exchanging rows $j$ and $\ell-j$ in $\mathbf{M}^{(\ell)}$ is equivalent to multiplying the argument of the symmetric polynomial in eq.˜9 by a factor of 1-1, we have
@@
where the second line follows from the identity $e_{k}(-\boldsymbol{v})=(-1)^{k}e_{k}(\boldsymbol{v})$. To show the result for the other ordering, we utilize the reciprocal polynomial of $p(x)$, given by $p^{*}(x)=x^{d}p(1/x)$, where $d$ is the degree of $p$. In particular, the degree-$k$ coefficient of $p$ is the degree-(dk)(d-k) coefficient of $p^{*}$. For any $k$, we have by lemma˜4
@@
Together, eq.˜100 and eq.˜105 complete the proof. ∎
B.3 Proof of lemma˜7
@@
Each factor in parentheses is a FLO transformation, since $\gamma_{j}\gamma_{k}=\exp\left(\frac{\pi}{2}\gamma_{j}\gamma_{k}\right)$, and $\gamma_{1}\in\mathds{F}(2n)$ by the definition below eq.˜3. Thus, we can perform the FLO twirl on $\mathcal{E}$ by first performing a Pauli twirl, mapping $\mathcal{E}\rightarrow\mathcal{E}^{\mathds{P}(n)}$, and then twirling over the full group of FLO unitary transformation, mapping $\mathcal{E}^{\mathds{P}(n)}\rightarrow(\mathcal{E}^{\mathds{P}(n)})^{\mathds{F}(2n)}=\mathcal{E}^{\mathds{F}(2n)}$.
By applying the Jordan-Wigner transform to eq.˜22, together with the first statement of the lemma, we have
$\mathcal{E}^{\mathds{F}(2n)}(X)$ $=\sum_{\boldsymbol{\alpha}\subseteq[2n]}p_{\mathbf{x}(\boldsymbol{\alpha})}\int_{\mathds{F}(2n)}dU\;\mathcal{U}\left(\gamma_{\boldsymbol{\alpha}}\right)X\mathcal{U}(\gamma^{\dagger}_{\boldsymbol{\alpha}})$ (108)
$\mathcal{E}^{\mathds{F}(2n)}(X)$ $=\sum_{\begin{subarray}{c}\boldsymbol{\alpha}\subseteq[2n]\\ \boldsymbol{\beta},\boldsymbol{\eta}||\boldsymbol{\alpha}|,2n\end{subarray}}p_{\mathbf{x}(\boldsymbol{\alpha})}\left[\int_{O(2n)}d\mathbf{R}\det\left(\mathbf{R}_{\boldsymbol{\alpha}\boldsymbol{\beta}}\right)\det\left(\mathbf{R}_{\boldsymbol{\alpha}\boldsymbol{\eta}}\right)\right]\left(\gamma_{\boldsymbol{\beta}}X\gamma_{\boldsymbol{\eta}}^{\dagger}\right)$ (109)
@@
$\mathcal{E}^{\mathds{F}(2n)}(X)$ $=\sum_{\begin{subarray}{c}\boldsymbol{\alpha}\subseteq[2n]\\ \boldsymbol{\beta}||\boldsymbol{\alpha}|,2n\end{subarray}}p_{\mathbf{x}(\boldsymbol{\alpha})}\left[\int_{O(2n)}d\mathbf{R}\ |\det\left(\mathbf{R}_{\boldsymbol{\alpha}\boldsymbol{\beta}}\right)|^{2}\right]\left(\gamma_{\boldsymbol{\beta}}\odot\gamma_{\boldsymbol{\beta}}^{\dagger}\right)$ (110)
@@
$\int_{O(2n)}d\mathbf{R}\ |\det\left(\mathbf{R}_{\boldsymbol{\alpha}\boldsymbol{\beta}}\right)|^{2}$ $=\int_{O(2n)}d\mathbf{R}\ |\det\left[\left(\mathbf{S}^{\mathrm{T}}\mathbf{R}\right)_{\boldsymbol{\alpha}\boldsymbol{\beta}}\right]|^{2}$ (111)
$\int_{O(2n)}d\mathbf{R}\ |\det\left(\mathbf{R}_{\boldsymbol{\alpha}\boldsymbol{\beta}}\right)|^{2}$ $=\int_{O(2n)}d\mathbf{R}\ |\det\left(\mathbf{R}_{\boldsymbol{\eta}\boldsymbol{\beta}}\right)|^{2}$ (112)
@@
$\int_{O(2n)}d\mathbf{R}\ |\det\left(\mathbf{R}_{\boldsymbol{\alpha}\boldsymbol{\beta}}\right)|^{2}$ $=\binom{2n}{|\boldsymbol{\alpha}|}^{-1}\sum_{\boldsymbol{\eta}||\boldsymbol{\alpha}|}\int_{O(2n)}d\mathbf{R}\ |\det\left(\mathbf{R}_{\boldsymbol{\eta}\boldsymbol{\beta}}\right)|^{2}$ (113)
$=\binom{2n}{|\boldsymbol{\alpha}|}^{-1}\int_{O(2n)}d\mathbf{R}\ \det\left[\left(\mathbf{R}^{\mathrm{T}}\mathbf{R}\right)_{\boldsymbol{\beta}\boldsymbol{\beta}}\right]$ (114)
$\int_{O(2n)}d\mathbf{R}\ |\det\left(\mathbf{R}_{\boldsymbol{\alpha}\boldsymbol{\beta}}\right)|^{2}$ $=\binom{2n}{|\boldsymbol{\alpha}|}^{-1}$ (115)
@@ B.4 Proof of lemma˜11
For an ensemble of noisy and FLO-twirled FLO circuits $\mathcal{C}^{\mathds{F}(2n)}_{r}$,
@@
We begin with the statement for $\mathcal{C}_{z}^{\mathds{F}(2n)}$. We have
@@
Applying the argument leading to eq.˜37 gives
@@
Circuits in $\mathcal{C}_{z}^{\mathds{F}(2n)}$ are chosen to interact as the identity on net. This gives
@@
$\mathrm{P}^{(j)}_{0}=2^{-n}\mathbf{d}^{(n)}\cdot\mathbf{M}^{(n)}\cdot\boldsymbol{\Lambda}_{\mathrm{even}}^{(j)}.$ (133)
In the $\mathcal{C}_{x}^{\mathds{F}(2n)}$-case, instead of eq.˜122, we have
@@
The corresponding probability for the “-” outcome on the first bit is given by
@@
$\frac{1}{2}(\mathrm{P}^{(j)}_{+}+\mathrm{P}^{(j)}_{-})$ $=2^{-n}\mathbf{d}^{(n-1)}\cdot\mathbf{M}^{(n-1)}\cdot\boldsymbol{\Lambda}_{\mathrm{even}\setminus\{2n\}}^{(j)}$ (148)
$\frac{1}{2}(\mathrm{P}^{(j)}_{+}-\mathrm{P}^{(j)}_{-})$ $=2^{-n}\mathbf{d}^{(n-1)}\cdot\mathbf{M}^{(n-1)}\cdot\boldsymbol{\Lambda}_{\mathrm{odd}}^{(j)}$ (149)